Please show all work

Ex: convert a) 670 ng to Mg b) c) d) e) 78nm3 to L Convert 864mg/CL to μg/ML 674 mm3 to gallons (1 gal=3.785L) Convert 467 Tm to miles. 1 mi-1.609 km.
Answer #1

(a) 670 ng = 670 ng x ( 10-9g/nm) x (10-3 kg/g) x ( 10-3 Mg/kg) = 670x10-15 Mg

(b) 78 nm3 = 78nm3x(10-9 m/nm) 3 x ( 103 L/m3) = 78x10-24 L

(c) 864 ng/cL = 864 ng/cL x ( 10-3\mug/ng) x (1 cL/10-8 ML) = 864x105\mug/ML

(d) 674 mm3 = 674 mm3 x ( 10-6 L/mm3) x (1 gal / 3.785 L) = 178.1x10-6 gallons

(e) 467 Tm = 467 Tm x ( 1012 m/Tm) x (10-3 km / m ) x ( 1 mi /1.609 km) = 2.90x1011 miles